# TLControl: Trajectory and Language Control for Human Motion Synthesis

### v3

@@ Abstract
Controllable human motion synthesis is essential for applications in AR/VR, gaming, movies, and embodied AI. Existing methods often focus solely on either language or full trajectory control, lacking precision in synthesizing motions aligned with user-specified trajectories, especially for multi-joint control. To address these issues, we present TLControl, a new method for realistic human motion synthesis, incorporating both low-level trajectory and high-level language semantics controls. Specifically, we first train a VQ-VAE to learn a compact latent motion space organized by body parts. We then propose a Masked Trajectories Transformer to make coarse initial predictions of full trajectories of joints based on the learned latent motion space, with user-specified partial trajectories and text descriptions as conditioning. Finally, we introduce an efficient test-time optimization to refine these coarse predictions for accurate trajectory control. Experiments demonstrate that TLControl outperforms the state-of-the-art in trajectory accuracy and time efficiency, making it practical for interactive and high-quality animation generation. Project Website: https://tlcontrol.weilinwl.com
Figure 1: TLControl, a novel method for Trajectory and Language Control for Human Motion Synthesis. The dotted lines represent the input control trajectories. The corresponding control joints are highlighted in orange. Our method demonstrates versatile multi-joint controls (see Figures 1a to 1c), the ability to handle complex trajectories (see Figure 1d), multi-stage control (see Figure 1e), and the preservation of language semantics while utilizing trajectory controls (see Figure 1f). We highly encourage readers to view our supplementary video to see our results.
@@ 1 Introduction
Human motion generation is a fundamental problem in computer graphics, with a wide range of applications in AR/VR, gaming industry and large-scale training data generation for embodied AI. In these applications, it is crucial to support user control at different levels on motion generation. For example, users may want to have high-level semantics control (e.g., text descriptions) for motion synthesis, and, meanwhile, expect flexible and precise spatial control to “drag” any joint on arbitrary trajectories. In this work, we study the problem of realistic and precise human motion generation with both high-level language and low-level trajectory controls.
@@
Recently, great advancements have been made in language-conditioned human motion generation [12, 41, 32, 13, 1, 8, 30, 48, 59, 23, 22, 39, 2]. However, most existing works only support language control [48, 7, 58]. Another line of research explores trajectory-controlled motion generation [38, 40, 5], a.k.a., inverse kinematics, but it lacks high-level language semantics control. Furthermore, inverse kinematics can only infer joint configurations from full (complete) trajectories of end-effectors, while our method generates full motions from partial (incomplete) trajectories. The most related works are PriorMDM[39], GMD [22], and a concurrent work, Omnicontrol [54], that incorporate spatial control signals in a language-conditioned motion generation model, but neither of them can accurately align the motion with specified trajectories, especially when controlling multiple joints simultaneously, as depicted in Figure 4. Moreover, their motion generation process is time-consuming due to inefficient multi-step sampling in the motion diffusion model, restricting its practicality in real-world applications.
@@
To address these issues, we propose TLControl for Trajectory and Language-controlled human body motion generation. We propose to operate in a part-structured latent encoding of human motions, learned with a VQ-VAE exploiting knowledge of the skeletal topology of human body. We then synthesize full-body motions from partial trajectory and language input in two steps: a feed-forward coarse estimation, and a subsequent optimization-based refinement. To efficiently train the feed-forward model, we propose a Masked Trajectory Transformer (MTT) that is trained by synthesizing various possible partial trajectory controls through spatial and temporal masking, enabling highly flexible operation. The optimization step starts at the MTT solution, and then more precisely aligns it with user-specified trajectories through an iterative search procedure.
@@
We conduct extensive experiments to validate the effectiveness of our approach. Compared to existing methods, TLControl exhibits superior performance in both trajectory accuracy and time efficiency. The results further illustrate that our method empowers users to interactively generate and modify high-quality animations within a brief runtime, affirming its practicality and effectiveness.
@@
    We introduce a new method TLControl that generates high-quality and precise human motions with highly flexible language and partial trajectory-based user specifications.
@@
    TLControl learns a human body morphology-aware structured latent space suitable for flexible motion synthesis.
@@
    An efficient feed-forward transformer model trained using carefully designed masking strategies initializes iterative optimization for large computational efficiencies in motion synthesis compared to prior work, at comparable or superior performance levels.
@@ 2 Related Work
There have recently been significant advances in human motion generation, incorporating diverse, multi-modal inputs such as text prompts [12, 41, 32, 48, 13, 1, 23, 29, 8, 30, 21], action labels [31, 10], partial motion sequences [9, 15, 51], control signals [45, 43, 28, 8], music [25, 24, 2], and images [38, 6]. In this section, we conduct a comprehensive review of the existing literature in the field of motion generation. Our reviews covers the studies ranging from unconditional motion generation to conditional approaches, particularly focusing on language-conditioned motion synthesis, known as action-to-motion and text-to-motion. We also explore the integration of trajectory or path controls in these processes, highlighting how these innovations contribute to advancing the field.
@@
Unconditional Motion Generation This area of research focuses on the autonomous generation of motion sequences without specific guiding conditions or annotations. Pioneering studies, such as those referenced in  [55, 61, 60, 36, 48] have made notable strides in this domain. These works stand out for their unique ability to capture and model the entire motion space, leveraging raw motion data to produce diverse and dynamic motion patterns. These methods exhibit innovation in interpreting and representing complex motion dynamics. However, as our research primarily focuses on conditional motion generation, our discussion of unconditional motion generation only serves as a contextual background to our primary focus.
@@
Conditional Motion Generation In the realm of conditional, tasks, ACTOR [31] presents a class-agnostic transformer VAE as a baseline. It introduces learnable biases within a transformer VAE to encapsulate action for motion generation. But similar to other works in the field of action-to-motion [10], it is limited to cover changes in the conditional inputs. Recent studies in the task text-to-motion [32, 1, 59, 48, 23, 12, 33, 3] has emerged as a principal driver, reshaping research frontiers with its user-friendly nature by allowing nature languages input. The advancements can be categorized into two methodologies: joint-latent models and diffusion models. Joint-latent models like TEMOS [32] typically employs a motion VAE alongside a text variational encoding. These components are then aligned into a compatible latent space by using divergence loss. This method advances the synthesis of human motion sequences from natural language inputs, representing a significant step forward in the field. However, as observed in the misalignment of Gaussian distributions [47], there are difficulties remained in aligning the structure and distribution of natural language and human motions.
@@
Recent advancements in image synthesis have been marked by the successful application of diffusion generative models [42]. Building on this progress, these models have now been extended to the domain of human motion synthesis. This novel application is being applied in recent works such as those presented in [48, 7, 19, 23, 54, 2]. Notably, MotionDiffuse [59] emerges as the first text-based motion diffusion model, offering fine-grained instructions for individual body parts. MDM [48] introduces a motion diffusion model that operates directly on raw motion data, further expanding the scope of natural language controls in human motion generation. This method represents a significant leap forward in terms of enabling intuitive, natural language inputs to motion synthesis.
@@ 3 Method
We aim to generate full-body motions to match user specifications, which consist of two parts: a text description $\mathbf{L}$, and a partial trajectory $\mathbf{R^{\prime}}$. $\mathbf{R^{\prime}}$ has $T$ frames, corresponding to the target motion length. In each frame $t$, $\mathbf{R^{\prime}}$ contains the 3D positions of some subset $m_{t}<n$ of the $n$ full-body joints, corresponding to the “control joints” which the user chose to specify for that frame. For example, the user might choose to specify only the head joints for all frames $t$, or even specify different joints in different frames. The desired final output is the full body motion $\mathbf{J}\in\mathbb{R}^{T\times M}$ of the character, where $M$ is the dimension of the single frame pose represented by the joint rotations and positions.
We propose to train such a system, TLControl, using data in which each sample $(\mathbf{R},\mathbf{J},\mathbf{L})$ is a tuple containing a full trajectory specifying some key joints, ground truth full body motions, and language description. To achieve the goal, we first learn a motion embedding by training a VQ-VAE [49] to establish part-based latent spaces, representing human body motion at the body part level (Sec.3.1). We then train a text-conditioned Masked Trajectory Transformer (MTT) to output a coarse completion of partially masked trajectories in this learned part-based embedding (Sec.3.2). This coarse completion is then further refined by a simple yet effective optimization over the learned latent space to minimize the distance between control targets and completed motions, leading to accurate and efficient trajectory control (Sec. 3.3). An overview of our framework is shown in Fig. 2.
Figure 2: Overview of TLControl framework: At training stage I, we train the part-based VQ-VAE in 3.1 for reconstructing human motions. In training stage II, the decoder of the part-based VQ-VAE is frozen and we train the masked trajectory transformer (MTT) in 3.2 for predicting code indices from control inputs. Finally, at test time, the MTT receives text description and partial control trajectories to predict an initial VQ-VAE quantized code seed, which is refined by run-time optimization as in 3.3 before decoding with the VQ-VAE into full body motions.
@@ 3.1 Part-based VQ-VAE
Learning discrete representations has proven effective in motion synthesis tasks in recent works [14, 58]. These methods use VQ-VAE [49] for discretizing a continuous motion space, facilitating the capturing of complex motion patterns in a structured manner and enhancing the generation of coherent and high-fidelity motion sequences. However, previous works [14, 58] encode the full body into the discrete motion space, treating the human body as a holistic entity. Instead, in this paper, we propose to learn a better representation of part-based motion priors, which have demonstrated impressive results for computer character animation [44, 4, 18, 34]. To do so, we propose a part-based VQ-VAE that disaggregates the motion at the human body-part level for learning a well-structured and compact latent space.
@@
where $\beta$ represents a weight hyper-parameter, while sg denotes the stop-gradient operator. Also, suggested by previous works in this field [58], we apply code reset and the exponential moving average [46] to prevent codebook collapse during training.
@@ 3.2 Masked Trajectory Transformer
Specifically, our Masked Trajectory Transformer (MTT) takes CLIP embeddings of the text prompts, as in prior work [58, 48]. For trajectories, we adopt the same joint partition in 3.1 and treat the trajectories of the five end-effectors and the root joint as the ground truth “full” trajectories $\mathbf{R}$ for training. Recall that in our application settings, we would like to enable users to provide rough partial sketches of the desired trajectories to control motion generation. To simulate such specifications when training the MTT, we apply two masking strategies to $\mathbf{R}$ to obtain the masked control trajectories $\mathbf{R^{\prime}}$:
@@
$\triangleright$ Joint-Level Masking. Recognizing that users may not always provide trajectories for all six joints, we include joint-level masking. This approach simulates scenarios where only a subset of joint trajectories are specified. Specifically, we randomly select a number of joints and mask all corresponding trajectory points during training.
@@
We train the MTT to process the user-provided observations $(\mathbf{R}^{\prime},\bm{L})$ to infer quantized latent VQ-VAE motion codes corresponding to the ground truth motions $\mathbf{J}$. Specifically, the MTT infers indices $I$. Indexing with $I$ into the part codebooks $\{C_{k}\}$ produces the concatenated quantized latent $\hat{\mathbf{Q}}_{0}$, which may be decoded using the frozen VQ-VAE decoder $\bm{\mathcal{D}}$. Thus, the loss function for training the MTT is:
@@
Please refer to our Suppl. Materials for more details.
@@ 3.3 Test-Time Optimization
Since our body-part level motion embedding effectively captures the motion distribution in a compact space, this simple yet effective optimization framework produces high-quality human motion with the desired controllability while maintaining semantic coherence with the textual description. In contrast to the closest prior work OmniControl [54] where limited control specifications are coupled with the network during training, this framework allows for flexible controls, such as simultaneously controlling multiple joints at different time steps; See Sec. 4.1. Moreover, it is much more computationally efficient compared with SOTA methods, which we investigate in Sec. 4.2.
@@ 4 Experiment
We conduct extensive experiments of our method and compare it with state-of-the-art methods in terms of motion quality, controllability, accuracy, and efficiency. We further investigate the key design choices by conducting ablation studies.

### v4

# denotes equal contributions

@@ Abstract
Controllable human motion synthesis is essential for applications in AR/VR, gaming and embodied AI. Existing methods often focus solely on either language or full trajectory control, lacking precision in synthesizing motions aligned with user-specified trajectories, especially for multi-joint control. To address these issues, we present TLControl, a novel method for realistic human motion synthesis, incorporating both low-level Trajectory and high-level Language semantics controls, through the integration of neural-based and optimization-based techniques. Specifically, we begin with training a VQ-VAE for a compact and well-structured latent motion space organized by body parts. We then propose a Masked Trajectories Transformer (MTT) for predicting a motion distribution conditioned on language and trajectory. Once trained, we use MTT to sample initial motion predictions given user-specified partial trajectories and text descriptions as conditioning. Finally, we introduce a test-time optimization to refine these coarse predictions for precise trajectory control, which offers flexibility by allowing users to specify various optimization goals and ensures high runtime efficiency. Comprehensive experiments show that TLControl significantly outperforms the state-of-the-art in trajectory accuracy and time efficiency, making it practical for interactive and high-quality animation generation.
Keywords:
Motion generation Text-to-motion Trajectory-to-motion
Figure 1: TLControl, a novel method for Trajectory and Language Control for Human Motion Synthesis. The corresponding control joints are highlighted in orange. Our method demonstrates versatile multi-joint controls (see Figures 1a to 1c), the ability to handle complex trajectories (see Figure 1d), multi-stage control (see Figure 1e), and the preservation of language semantics while utilizing trajectory controls (see Figure 1f). The dotted lines represent the input control trajectories defined by users through specifying the parameters of analytical shapes. Note the input trajectories can also be hand drawings from users, or parameters from environment settings (See Fig. 4). We highly encourage readers to view our supplementary video to see our results.
@@ 1 Introduction
Human motion generation is a fundamental problem in computer vision, robotics and graphics, with diverse applications in augmented and virtual reality, the gaming industry, and embodied. In these applications, it is crucial to support user control at different levels on motion generation. For example, users may want to have high-level semantics control (e.g., text descriptions) for motion synthesis, and, meanwhile, expect flexible and precise spatial control to “drag" any joint on arbitrary trajectories. In this work, we study the problem of realistic and precise human motion generation with both high-level language and low-level trajectory controls; See an overview in Figure 1.
@@
Recently, great advancements have been made in language-conditioned human motion generation [14, 47, 38, 16, 1, 12, 35, 56, 70, 27, 26, 45, 2, 10]. However, most existing works only support language control [56, 6, 67]. Another line of research explores trajectory-controlled motion generation [44, 46, 4], a.k.a., inverse kinematics, but it lacks high-level language semantics control. Furthermore, inverse kinematics can only infer joint configurations from full (complete) trajectories of end-effectors, while our method generates full motions from partial (incomplete) trajectories. The most related works are PriorMDM[45], GMD [26], and a concurrent work, Omnicontrol [62], that incorporate spatial control signals in a language-conditioned motion generation model, but neither of them can accurately align the motion with specified trajectories, especially when controlling multiple joints simultaneously, as evidenced by Figure 8 and Figure 8. Moreover, their motion generation process is time-consuming (see Sec. 4.2) due to inefficient multi-step sampling in the motion diffusion model, restricting its practicality in real-world applications.
@@
To address these challenges, we propose TLControl, leveraging an integrated approach combining neural-based and optimization-based methods, for Trajectory and Language-controlled human body motion generation. Specifically, we propose to operate in a part-structured latent encoding of human motions, learned with a VQ-VAE exploiting knowledge of the skeletal topology of human body. We then synthesize full-body motions from partial trajectory and language input in two steps: a feed-forward coarse estimation, and a subsequent optimization-based refinement. To efficiently sample the feed-forward coarse estimations, we propose a Masked Trajectory Transformer (MTT) that predicts a distribution conditioned on trajectory controls and text descriptions. Then we present an optimization scheme begins with the MTT solution and then iteratively refines it to align with user-specified trajectories. Compared with previous methods, our framework offers flexible controls by allowing users to specify various optimization goals (Sec. 4.1) while enjoying high efficiency compared with SOTA methods (Sec. 4.2).
@@
As a result, TLControl produces high-quality motion that aligns with high-level language semantics while accurately following low-level control trajectories derived from 1) specifying parameters of analytical shapes, 2) hand drawings in 3D space, and 3) environment settings. This versatility highlights the practical applicability of our method in a wide range of scenarios. We conduct extensive experiments to validate the effectiveness of our approach. Compared to existing methods, TLControl exhibits superior performance in both trajectory accuracy and time efficiency. The results further illustrate that our method empowers users to interactively generate and modify high-quality animations within a brief runtime, affirming its practicality and effectiveness.
@@
    We introduce a novel method TLControl that generates high-quality and precise human motions with language and partial trajectory-based user specifications through the tight collaboration of neural-based and optimization-based methods.
@@
    Our framework effectively learns a compact human body morphology-aware structured latent space using a VQ-VAE, and a motion distribution conditioned on language and trajectory using a Masked Trajectory Transformer. The proposed optimization scheme offers flexible controls by allowing users to specify various optimization goals while ensuring high runtime efficiency.
@@
    TLControl demonstrates its superior control accuracy in producing realistic human motion with remarkable runtime efficiency, compared with SOTAs. Our codes will be publicly available upon the paper’s publication.
@@ 2 Related Work
There have recently been significant advances in human motion generation, given its wide applications across virtual reality [61, 28, 66], game development [54, 51, 52, 21, 50], human behavior analysis [33, 18, 72, 71, 65, 8, 11] and robotics [59, 48, 31, 63, 9]. Conditional motion generation incorporates diverse, multi-modal inputs such as text prompts [14, 47, 38, 56, 16, 1, 27, 68], action labels [37, 25, 12, 17], partial motion sequences [13, 19, 60], control signals [50, 51, 36, 12, 69, 5], music [30, 29, 2], and images [44, 7]. In this section, we conduct a comprehensive review of the existing literature in the field of motion generation. Our reviews covers the studies ranging from unconditional motion generation to conditional approaches, particularly focusing on language-conditioned motion synthesis, known as action-to-motion and text-to-motion. We also explore the integration of trajectory or path controls in these processes, highlighting how these innovations contribute to advancing the field.
@@
Unconditional Motion Generation This area of research focuses on the autonomous generation of motion sequences without specific guiding conditions or annotations. Pioneering studies, such as those referenced in [64, 74, 73, 42, 56] have made notable strides in this domain. These works stand out for their unique ability to capture and model the entire motion space, leveraging raw motion data to produce diverse and dynamic motion patterns. These methods exhibit innovation in interpreting and representing complex motion dynamics. As we primarily focus on conditional motion generation, our discussion of unconditional motion generation only serves as a contextual background to our primary focus.
@@
Conditional Motion Generation In the realm of conditional, tasks, ACTOR [37] presents a class-agnostic transformer VAE as a baseline. It introduces learnable biases within a transformer VAE to encapsulate action for motion generation. But similar to other works in the field of action-to-motion [17], it is limited to cover changes in the conditional inputs. Recent studies in the task text-to-motion [38, 1, 70, 56, 27, 14, 39, 3] has emerged as a principal driver, reshaping research frontiers with its user-friendly nature by allowing nature languages input. The advancements can be categorized into two methodologies: joint-latent models and diffusion models. Joint-latent models like TEMOS [38] typically employs a motion VAE alongside a text variational encoding. These components are then aligned into a compatible latent space by using divergence loss. This method advances the synthesis of human motion sequences from natural language inputs, representing a significant step forward in the field. However, as observed in the misalignment of Gaussian distributions [55], there are difficulties remaining in aligning the joint distribution of language and human motions.
@@
Recent advancements in image synthesis have been marked by the successful application of diffusion generative models [49]. Building on this progress, these models have now been extended to the domain of human motion synthesis. This novel application is being applied in recent works such as those presented in [56, 6, 24, 27, 62, 58, 2]. Notably, MotionDiffuse [70] emerges as the first text-based motion diffusion model, offering fine-grained instructions for individual body parts. MDM [56] introduces a motion diffusion model that operates directly on raw motion data, further expanding the scope of natural language controls in human motion generation. This method represents a significant leap forward in terms of enabling intuitive, natural language inputs to motion synthesis.
@@ 3 Method
We aim to generate full-body motions to match user specifications, which consist of two parts: a text description $\mathbf{L}$, and a partial trajectory $\mathbf{R^{\prime}}$. $\mathbf{R^{\prime}}$ has $T$ frames, corresponding to the target motion length. Notably, we assume that there is no conflict between language and the trajectory lines, and the trajectories themselves represent motions that are feasible in reality. In each frame $t$, $\mathbf{R^{\prime}}$ contains the 3D positions of some subset $m_{t}<n$ of the $n$ full-body joints, corresponding to the “control joints” which the user chose to specify for that frame. For example, the user might choose to specify only the head joints for all frames $t$, or even specify different joints in different frames. The desired final output is the full body motion $\mathbf{J}\in\mathbb{R}^{T\times M}$ of the character, where $M$ is the dimension of the single frame pose represented by the joint rotations and positions following [15].
We train TLControl using data in which each sample $(\mathbf{R},\mathbf{J},\mathbf{L})$ is a tuple containing a full trajectory specifying some key joints, ground truth full body motions, and language description. To achieve the goal, we first learn a motion embedding by training a VQ-VAE [57] to establish part-based latent spaces, representing human body motion at the body part level (Sec.3.1). We then train a text-conditioned Masked Trajectory Transformer (MTT) to output a coarse completion of partially masked trajectories in this learned part-based embedding (Sec.3.2). This coarse completion is then further refined by a simple yet effective optimization over the learned latent space to minimize the distance between control targets and completed motions, leading to accurate and efficient trajectory control (Sec. 3.3). An overview of our framework is shown in Fig. 2.
Figure 2: Overview of TLControl framework: At training stage I, we train the part-based VQ-VAE in 3.1 for reconstructing human motions. In training stage II, the decoder of the part-based VQ-VAE is frozen and we train the masked trajectory transformer (MTT) in 3.2 for predicting code indices from control inputs. Finally, at test time, the MTT receives text description and partial control trajectories to predict an initial VQ-VAE quantized code seed, which is refined by run-time optimization as in 3.3 before decoding with the VQ-VAE into full body motions.remaking.
@@ 3.1 Part-based VQ-VAE
Learning discrete representations has proven effective in motion synthesis tasks in recent works [16, 67]. These methods use VQ-VAE [57] for discretizing a continuous motion space, facilitating the capturing of complex motion patterns in a structured manner and enhancing the generation of coherent and high-fidelity motion sequences. However, previous works [16, 67] encode the full body into the discrete motion space, treating the human body as a holistic entity. Instead, in this paper, we propose to learn a better representation of part-based motion priors, which have demonstrated impressive results for computer character animation [22, 40, 52]. To do so, we propose a part-based VQ-VAE that disaggregates the motion at the human body-part level for learning a well-structured and compact latent space.
@@
where $\beta$ is a balancing term, while sg[$\cdot$] denotes the stop-gradient operator. Also, suggested by previous works in this field [67], we apply code reset and the exponential moving average [53] to prevent codebook collapse during training.
@@ 3.2 Masked Trajectory Transformer
Specifically, our Masked Trajectory Transformer (MTT) takes CLIP embeddings of the text prompts, as in prior work [67, 56]. For trajectories, we adopt the same joint partition in Sec. 3.1 and treat the trajectories of the five end-effectors and the root joint as the ground truth “full” trajectories $\mathbf{R}$ for training. Recall that in our application settings, we would like to enable users to provide rough partial sketches of the desired trajectories to control motion generation. To simulate such specifications when training the MTT, we apply two masking strategies to $\mathbf{R}$ to obtain the masked control trajectories $\mathbf{R^{\prime}}$:
@@
$\triangleright$ Joint-Level Masking. Since users may not always provide trajectories for all six joints, we include joint-level masking. This simulates scenarios where only a subset of joint trajectories are specified. Specifically, we randomly select a number of joints and mask corresponding trajectory points during training.
@@
Sampling the initial quantized latent $\hat{\mathbf{Q}}_{0}$ requires training the MTT to process the user-provided observations $(\mathbf{R}^{\prime},\bm{L})$ and outputs the logits of selecting code indices $I$. We transform the logits of each code into a one-hot decision vector $\mathbf{v}$ using the Gumbel-Softmax [23], a method that enables sampling from distributions by integrating Gumbel noise. In other words, the sampling capability allows for generative motion synthesis. Then, each code is selected using $\mathbf{C}_{k}\cdot\mathbf{v}$, and the resultant selected codes are concatenated to form $\hat{\mathbf{Q}}_{0}$, which can be decoded using the frozen VQ-VAE decoder $\bm{\mathcal{D}}$. Thus, the loss function for training the MTT is:
@@ 3.3 Test-Time Optimization
Since our body-part level motion embedding effectively captures the motion distribution in a compact space, this simple yet effective optimization framework produces high-quality human motion with the desired controllability while maintaining semantic coherence with the textual description. In contrast to the closest prior work OmniControl [62] where limited control specifications are coupled with the network during training, this framework allows for flexible controls, such as simultaneously controlling multiple joints at different time steps. Moreover, it is much more computationally efficient compared with SOTA methods, which we investigate in Sec. 4.2.
@@ 4 Experiment
We conduct extensive experiments of our method and compare it with state-of-the-art methods in terms of motion quality, controllability, accuracy, and efficiency. We further investigate the key design choices in ablation studies.
Figure 3: Qualitative results of TLControl using user-defined trajectories. Figure 3a and Figure 3d demonstrate that our method enables separate controls using language and joint-level trajectories. Figure 3b and Figure 3c showcase the capability of our method to manage multi-joint control simultaneously. Please refer to our supplementary for more qualitative results
Figure 4: Figure 4a to 4d: Qualitative results of our method using 3D user hand drawings on the left. Figure 4e and Figure 4f: Showcase that allocating foot placements according to the staircases.
4.0.1 Datasets
4.0.2 Metrics
Following the experiments setting in [62], we evaluate the fidelity of our result using Frechet Inception Distance (FID), R-Precision and Diversity. FID measures the quality of motion created by generative techniques. It calculates the disparity between actual and produced distributions within the feature space of a pre-trained model as being used in [15]. R-Precision measures a motion sequence against 32 text narratives, one of which is accurate and the remaining 31 are random descriptions. It involves computing the Euclidean distances between the embeddings of the motion and the textual narratives. The accuracy of retrieving the correct text from the motion is then reported based on the top-3 matches. Diversity is evaluated by randomly pairing all generated sequences from the test texts. The average cumulative difference within each pair is then calculated to determine diversity. MModality is the average output variance given the same input.